A competitive firm's cost of producing q units of output is C = 18 + 4q + q^2 Its corresponding marginal cost is MC = 4 + 2q. a. The firm faces a market price p = $24. Create a spreadsheet with q = 0, 1, 2, ..... 15, where the columns are q, R, C, VC, AVC, MC, and profit. Determine the profit-maximizing output for the firm and the corresponding profit. Should the firm produce this level of output or should it shut down? Explain. b. Suppose the competitive price declines to p = $12. Repeat the calculations of part a. Should the firm shut down? Please show Excel formulas. Thanks!
Answer #1

(a)

R = p x q = $24 x q

C = 18 + 4q + q2

VC = 4q + q2

AVC = VC/q = 4 + q

MC = dC/dq = 4 + 2q

Profit = R - C

q R C VC AVC MC Profit
0 0 18 0 -18
1 24 23 5 5 5 1
2 48 30 12 6 7 18
3 72 39 21 7 9 33
4 96 50 32 8 11 46
5 120 63 45 9 13 57
6 144 78 60 10 15 66
7 168 95 77 11 17 73
8 192 114 96 12 19 78
9 216 135 117 13 21 81
10 240 158 140 14 23 82
11 264 183 165 15 25 81
12 288 210 192 16 27 78
13 312 239 221 17 29 73
14 336 270 252 18 31 66
15 360 303 285 19 33 57

Profit is maximized (= $82) when q = 10 units. At this output, AVC < Price ($14 < $24), so firm should continue to operate.

(b)

Price = $12
q R C VC AVC MC Profit
0 0 18 0 -18
1 12 23 5 5 5 -11
2 24 30 12 6 7 -6
3 36 39 21 7 9 -3
4 48 50 32 8 11 -2
5 60 63 45 9 13 -3
6 72 78 60 10 15 -6
7 84 95 77 11 17 -11
8 96 114 96 12 19 -18
9 108 135 117 13 21 -27
10 120 158 140 14 23 -38
11 132 183 165 15 25 -51
12 144 210 192 16 27 -66
13 156 239 221 17 29 -83
14 168 270 252 18 31 -102
15 180 303 285 19 33 -123

In this case firm is making loss at every output level, and loss is minimized (= -$2) when q = 4 units. At this output, AVC < Price ($8 < $12), so firm should not shut down.

E 520 PC Cost - Microsoft Excel Home Insert Page Layout Formulas Data Review View novaPOF e JFormat Painter B 1 u、ロー 3_▲▼ 븀 谭